# A Unified Framework for Iris Anti-Spoofing: Introducing IrisGeneral Dataset and Masked-MoE Method

### v1

@@ Introduction
Table 2: This table shows the Train / Test usage of each dataset of each protocol. In Protocol-1, all datasets’ training and testing sets are involved. In Protocol-2, training sets are H and F, and testing sets involve all the 10 subsets. In Protocol-3, for each sub-protocol, we choose one device to test, and others to train.
@@
Protocols of IrisGeneral
Three protocols are defined in the IrisGeneral dataset to evaluate the three kinds of generalization performance of the models. Table 2 shows the Train / Test usage of each dataset.
@@
Proposed Method
In this work, We proposed the Masked-MoE (MMoE) method with two main contributions: Mask Mechanism and Cosine Distance Loss. Figure 3 demonstrates the critical part of the framework, and Algorithm 1 provides the working process, including the details of masking dispatch weights and the measure of cosine distances between experts. The MMoE is based on SoftMoE (Puigcerver et al. 2023), and could easily added to the image encoder of CLIP (Radford et al. 2021) as shown in the Figure 4.
Mask Mechanism of MMoE
The overfitting problem of MoE methods inherits exists and is hard to solve (Aceña et al. 2022; Chang et al. 2013; Allingham et al. 2021; Xie et al. 2023). Inspired by MAE (He et al. 2022), which improves the generalization of auto-encoder architecture by masking patches of samples, we refer to this masking method to improve the generalization to deal with the overfitting of MoE. We proposed a novel approach Mask Mechanism, which is different from the sample-wise masking manner of MAE, it is a feature-wise Mask Mechanism that masks parts of the latent features of images by tokens. To achieve that, we generate a random Mask Matrix with fixed mask rates, then use it to mask the Dispatch Weights by Hadamard product. Then the input tokens are weighted by Masked Dispatch Weights, and thus parts of them are masked as shown in Figure 3.
@@
$\mathbf{X}\in\mathbb{R}^{{n\times d}}$ is the input tokens of MMoE, n denotes the number of tokens and d denotes their feature dimension. In MMoE with e experts, each expert deals with s slots where each slot has a corresponding vector of parameters with d-dimension. $\mathbf{\Phi}\in\mathbb{R}^{{d\times(e\cdot s)}}$ is a randomly initialized vector to weight the input tokens. $\mathbf{D}\in\mathbb{R}^{{n\times(e\cdot s)}}$ is the dispatch weights, which is the result of applying a softmax on the columns of $\mathbf{X}\mathbf{\Phi}$. $\mathbf{D}_{ij}$ denotes the element in dispatch weights $\mathbf{D}$ located in the ${i}^{th}$ column and ${j}^{th}$ row. $\tilde{\mathbf{X}}\in\mathbb{R}^{(e\cdot s)\times d}$ is the input of Experts, a set of slots. It is obtained by weighting $\mathbf{X}$ with the dispatch weights $\mathbf{D}$. The scale of our proposed Mask Matrix is the same scale as the one of Dispatch Weights $\mathbf{D}$, $\mathbf{M}\in\mathbb{R}^{{n\times(e\cdot s)}}$, which consists only of 0 and 1. The mask rate denotes the proportion of 0 in Mask Matrx. In practice, we employ 4 experts in each block of the backbone of CLIP’s image encoder. Only one expert could get the complete information of all slots, while others could only get the masked slots. And then our Masked Dispatch Weights $\mathbf{D^{\prime}}$ is the Hadamard product of Dispatch Weights $\mathbf{D}$ and Mask Matrix $\mathbf{M}$, as shown in Equation 2.
@@
$\mathbf{D^{\prime}}_{ij}$ and $\mathbf{M}_{ij}$ denotes the element in masked dispatch weights $\mathbf{D^{\prime}}$ and Mask Matrix $\mathbf{M}$ located in the ${i}^{th}$ column and ${j}^{th}$ row. $\mathbf{X}$ is then weighted by $\mathbf{D^{\prime}}$ and we get $\tilde{\mathbf{X}}$, the input of Experts.
@@
Cosine Distance Loss between Experts
The efficiency of MoE methods is mainly thanks to the sparse characteristically, which means that each expert could learn different knowledge (Cai et al. 2024). So if we could find a way to let experts learn the knowledge from others, it may improve the generalization of the models. To address that, based on the Mask Mechanism, we calculate the distance between the output of experts, to encourage each expert to guess the unseen part and make up the masked information. Due to the aim of MoE that choose different experts focusing knowledge, the Mean Squared Error (MSE) which measures the absolute error between two tensors, is not suitable for evaluating the similarity between the outputs of experts. Thus we choose the cosine distance to calculate the similarity between the output of each expert and not request the same. Specifically, we masked three experts, the remaining one could see all of the information and request three masked experts’ output similar to the remaining one. As shown in Algorithm 1, all of the outputs of the masked experts request to calculate the cosine distance with the unmasked one. Equation 3 denotes the calculation process.
@@
$\tilde{Y^{i}_{j}}$ denotes the $j^{th}$ slot of the $i^{th}$ Expert’s output. The Cosine Distance is the mean of the cosine similarity between the $j^{th}$ slots of the two Experts. The $e$ is the number of Experts and the $s$ is the number of slots. The $cosine\_distance_{k}$ denotes the $k^{th}$ residual block of ViT-B/16, which is the backbone of our image encoder. As shown in Equation 4, the $cosine\_distance$ is used to calculate our mmoe loss $L_{mmoe}$.
@@
Table 3: This table shows the evaluation results on Protocol-1, Protocol-2, and Protocol-3 (Average).
@@
Table 4: This table shows the specific evaluation results of each sub-protocol in Protocol-3.
Experiment
Datasets and Experimental Setting
@@
We chose our proposed IrisGeneral as the dataset, the details are shown in the Section IrisGeneral Dataset with Table 1. All three protocols of the IrisGeneral listed in Table 2 are used. Protocol-1 evaluates the average performance of models in the iris anti-spoofing task, and Protocol-2 and Protocol-3 evaluate the cross-racial adaption and cross-device adaption respectively. We choose the Average Classification Error Rate (ACER), the Overall Detection Accuracy (ACC), the Area Under the Curve (AUC), and the Equivalent Error Rate (EER) as performance evaluation metrics. The experimental details are as follows: The backbone of CLIP’s image encoder is ViT-B/16. The mask rate is set as ”0, 0.1, 0.1, 0.1.”, and the prompts are set as ”This is an image of <label>eye.” We employ NVIDIA A100 GPUs to train and test with the Adam optimizer and the learning rate is $10^{-6}$.
Figure 5: The feature distribution of each method on Protocol-1 of the IrisGeneral dataset. Our proposed method has the clearest border and regular shape.
Performance Evaluation
@@
The ResNet50 (He et al. 2016), ViT-B/16 (Dosovitskiy et al. 2020), CLIP (Radford et al. 2021), and FLIP (Srivatsan, Naseer, and Nandakumar 2023) are chosen as representative models, then benchmarked them with ACER, ACC, AUC, and EER. Table 3 shows all the evaluation results including our proposed MMoE with CLIP and other models on Protocol-1, Protocol-2, and Protocol-3 (Average). Table 4 provides the specific results of each sub-protocol in Protocol-3. The evaluation results show that our proposed MMoE method with CLIP gets the best performance on both three protocols of the IrisGeneral dataset. For Protocol-1, Ours achieves 1.23% ACER, which is a nearly 13% improvement compared to the vanilla CLIP with 1.41% ACER. It means that our MMoE method could effectively detect spoofing iris samples on all subsets of the IrisGeneral dataset and achieve SOTA performance. For Protocol-2, Ours achieves 7.59% ACER, much better than other previous models. It proves our proposed method has a high generalization ability on cross-racial challenges. Verified the effectiveness of our contributions. For Protocol-3, Ours achieves 4.85% ACER which has significant advantages over other previous models. It proves our proposed method also has a high generalization ability on cross-device challenges. In summary, above all the evaluations, our proposed MMoE method could effectively detect the spoofing iris images and perform high generalization on both cross-racial and cross-device situations. It achieves the SOTA performance on all protocols of the IrisGeneral dataset. Furthermore, Figure /reffig: distributions demonstrates the feature distribution of each method on Protocol-1. Our proposed method has the clearest border and regular shape in the latent space, showing its good ability in this iris anti-spoofing field.
Mask Rate Analysis
To discuss the impact of different rates of the mask, we set up 4 sets of mask rates and tested them on both Protocol-1 and Protocol-2. Among that, the first two sets mask the same rate for both three experts, and one expert could see all tokens with complete information. The third one masks different rates for three experts while one expert could also see all information as similar to the first two sets. The last one is quite different, in that all experts are masked part of tokens, which means the MMoE model could not gain the complete input. The details of each set are shown as follows:
Rate_1: 0, 0.1, 0.1, 0.1. Rate_2: 0, 0.5, 0.5, 0.5.
Rate_3: 0, 0.15, 0.3, 0.45. Rate_4: 0.1, 0.1, 0.1, 0.1.
According to the evaluation results shown in Table 5. The Rate_1 has the best performance on both Protocol-1 and Protocol-2, in which the ACERs are 1.23% and 7.59% respectively. The Rate_2 is a bit worse than the Rate_1 may caused by the high mask rate would make experts lose too much information and then influence the learning space of the MMoE model. The Rate_3 with different mask rates set has the worst performance on Protocol-1. This difference may increase the gap between the learning ability of experts and then cause a negative impact on the training process. The Rate_4 has a middle performance among these mask rate sets on Protocol-1, a bit worse than the Rate_1. And is the worst one on Protocol-2. This set makes the MMoE model could not gain the complete input information, which may influence the learning space.
@@
Table 5: The evaluation results of our MMoE with CLIP under different sets of mask-rates. The Rate_1 has the best performance on both Protocol-1 and Protocol-2.
Ablation Study

### v2

@@ II-C Cross-domain Adaption in Biometric Anti-Spoofing
Figure 3: This figure shows the Train / Test usage of each dataset of the IAS-CDT Protocol. In Protocol-1, all datasets’ training and testing sets are involved. In Protocol-2, training sets are H and F, and testing sets involve all 10 subsets. In Protocol-3, for each sub-protocol, we choose one device to test and others to train. And some image samples are shown, including 10 subsets with Real and Fake iris images.
II-D Mixture-of-Experts (MoEs)
The conceptual origins of MoE trace back to the work of Jacobs et al. [38], which proposed a modular architecture with competing expert networks for complex tasks like vowel recognition. This design decomposes tasks into subtasks, reducing interference. The system includes multiple feedforward networks (experts) and a gating network. Each expert receives the same input and produces outputs of the same size, while the gating network determines which expert to activate for each case. Compared to traditional single networks, this modular design and expert selection mechanism allow for better handling of complex tasks and improved generalization. This work is widely regarded as the seminal study of MoE.
Nowadays, MoE has regained attention due to its ability to handle complex tasks. The sparsity-driven conditional computation started from the work of Shazeer et al. [40]. They introduced a novel sparse gating mechanism for MoE layers, aiming to significantly increase the model capacity of neural networks through conditional computation, addressing the limitations of traditional models in handling large-scale data. Since the Transformer-based models have performed well on various tasks, Lepikhin et al. [41] were the first to apply the MoE structure to the Transformer architecture. This approach leverages a sparse activation MoE model with conditional computation, improving model quality without significantly increasing computational cost.
To further improve the stabilization of MoE, Zoph et al. [43] proposed the ST-MoE model, which provides a systematic framework for sparse expert model design through architectural principles and routing algorithms. The ST-MoE model employs a sparse expert architecture where each input is processed by only a small subset of experts, allowing for a significant increase in parameter count without a corresponding rise in computation. Through the introduction of sparse expert architecture and a suite of stability techniques, the ST-MoE model demonstrates the potential of sparse models in transfer learning tasks.
To address the challenges brought by the sparse structure of MoE, fully differential MoE appears. Puigcerver et al. [68] proposed a novel model, Soft MoE (Soft Mixture of Experts). Soft MoE assigns input tokens to expert slots using a weighted average computation, with each slot processed by a corresponding expert. By implementing a fully differentiable soft allocation mechanism, Soft MoE addresses common challenges in sparse MoE models, such as training instability, token dropout, poor scalability with expert numbers, and suboptimal fine-tuning performance, while maintaining the efficiency and capacity benefits of MoE.
III Iris Anti-Spoofing Cross-Domain-Testing Protocol
III-A Subsets Detail
There are 10 subsets, belonging to 7 databases, published by 4 institutions, collected with 6 types of devices in the Cross-Testing Protocol. Table I lists the information for each subset. Figure 3 provides some real and fake iris image samples of each subset.
CASIAH100 [15] with 6744 samples in total, collected by H100, belonging to database CASIAH100, contains 5398 samples in its training set and 1346 samples in its testing set. We call this dataset H in short.
Clarkson2013Dalsa [16] with 1356 samples in total, collected by DALSA, belonging to database Clarkson 2013 for LivDet-Iris 2013, contains 670 samples in its training set and 686 samples in its testing set. We call this dataset Da in short.
Clarkson2015Dalsa [16] with 2509 samples in total, collected by DALSA, belonging to database Clarkson 2015 for LivDet-Iris 2015, contains 1573 samples in its training set and 936 samples in its testing set. We call this dataset Db in short.
Clarkson2015LG2200 [16] with 1944 samples in total, collected by LG2200, belonging to database Clarkson 2015 for LivDet-Iris 2015, contains 990 samples in its training set and 954 samples in its testing set. We call this dataset La in short.
Clarkson2017LG2200 [16] with 5841 samples in total, collected by LG2200, belonging to the database Clarkson 2017 dataset for LivDet-Iris 2017, contains 3591 samples in its training set and 2250 samples in its testing set. We call this dataset Lb in short.
IFVEAI1000 [17] with 50000 samples in total, collected by AI1000, belonging to database IFVE, contains 40000 samples in its training set and 10000 samples in its testing set. And we call this dataset F in short.
ND2013LG4000 [18]with 4200 samples in total, collected by LG4000, belonging to database Notre Dame 2013 for LivDet-Iris 2013, contains 3000 samples in its training set and 1200 samples in its testing set. We call this dataset Ga in short.
ND2013AD100 [18]with 900 samples in total, collected by AD100sensor, belonging to database Notre Dame 2013 for LivDet-Iris 2013, contains 600 samples in its training set and 300 samples in its testing set. We call this dataset Aa in short.
ND2017LG4000 [16] with 3000 samples in total, collected by LG4000, belonging to database Notre Dame 2017 for LivDet-Iris 2017, contains 1200 samples in its training set and 1800 samples in its testing set. We call this dataset Gb in short.
ND2017-CDAD100 [16] with 1800 samples in total, collected by AD100sensor, belonging to database Notre Dame 2017 for LivDet-Iris 2017, contains 0 samples in its training set and 1800 samples in its testing set. We call this dataset Ab in short.
@@ III-A Subsets Detail
III-B Sub-Protocols of IAS-CDT Protocol
Three sub-protocols are defined in the IAS-CDT Protocol to evaluate the three kinds of generalization performance of the models. Figure 3 shows the Train / Test usage of each dataset of the IAS-CDT Protocol.
@@ III-B Sub-Protocols of IAS-CDT Protocol
IV Proposed Method
We propose this Masked-MoE anti-spoofing method (shorten as MMoE in the following) to improve the capability of iris recognition system on differentiating live irises and irises with contact lenses. The proportion of texture differences between live irises and those with contact lenses is relatively small in iris anti-spoofing tasks, which cause the difficulties on distinguishing them in practical application scenes. The proposed MMoE contains two main components, mask mechanism and cosine distance loss, which could improve the generalization and reduce the overfitting problem.
Because of the variety of iris collection devices, the iris recognition models would perform weak on cross-device scenes. To improve the cross-domain capability of anti-spoofing model, we introduce MoE to alleviate the negative impacts caused by cross-domain datasets. While MoE improves model generalization, the increased parameters combined with limited discriminative features may lead to overfitting. To address this, a masking mechanism is introduced to guide experts in focusing on critical features for anti-spoofing decisions by dropping out redundant information. And then, to mitigate gaps between experts handling masked and unmasked features, comparisons are made between their outputs, leveraging complementarity among experts to enhance overall model generalization.
Figure 4 demonstrates the critical part of the framework, and Algorithm 1 provides the working process, including the details of masking dispatch weights and the measure of cosine distances between experts. Compared with soft MoE [68], our MMoE masked parts of the features before putting them into experts and calculating the cosine distances of the output of different experts. This MMoE is easily to be added to the image encoder of CLIP [33] as shown in the Figure 5.
IV-A Mask Mechanism of MMoE
As mentioned before, MoE could improve generalization but risks overfitting due to increased parameters and insufficient discriminative features. This overfitting problem of MoE methods inherits exists and is hard to solve [69, 70, 71, 72]. Inspired by MAE [50], which improves the generalization of auto-encoder architecture by masking patches of samples, we designed a masking mechanism where different experts focus on masked portions of features, guiding the model to learn key anti-spoofing characteristics. This novel approach Mask Mechanism, which is different from the sample-wise masking manner of MAE, it is a feature-wise Mask Mechanism that masks parts of the latent features of images by tokens. To achieve that, we generate a random Mask Matrix with fixed mask rates, then use it to mask the Dispatch Weights by Hadamard product. Then the input tokens are weighted by Masked Dispatch Weights, and thus parts of them are masked as shown in Figure 4. Moreover, considering the limited and subtle differences between live irises and those with contact lenses, the mask rate ought to be in a reasonable range, which could not only reduce the overfitting but also avoid omitting the obvious difference between liven irises and the fake ones.
@@ IV-A Mask Mechanism of MMoE
This Dispatch Weights is the original one in Soft MoE. $\mathbf{X}\in\mathbb{R}^{{n\times d}}$ is the input tokens of MMoE, n denotes the number of tokens and d denotes their feature dimension. In MMoE with e experts, each expert deals with s slots where each slot has a corresponding vector of parameters with d-dimension. $\mathbf{\Phi}\in\mathbb{R}^{{d\times(e\cdot s)}}$ is a randomly initialized vector to weight the input tokens. $\mathbf{D}\in\mathbb{R}^{{n\times(e\cdot s)}}$ is the dispatch weights, which is the result of applying a softmax on the columns of $\mathbf{X}\mathbf{\Phi}$. $\mathbf{D}_{ij}$ denotes the element in dispatch weights $\mathbf{D}$ located in the ${i}^{th}$ column and ${j}^{th}$ row. $\tilde{\mathbf{X}}\in\mathbb{R}^{(e\cdot s)\times d}$ is the input of Experts, a set of slots. It is obtained by weighting $\mathbf{X}$ with the dispatch weights $\mathbf{D}$.
Equation 3 denotes the main contribution of our Mask Mechanism. It shows the computation process of masking the dispatch weights. The scale of our proposed Mask Matrix is the same scale as the one of Dispatch Weights $\mathbf{D}$, $\mathbf{M}\in\mathbb{R}^{{n\times(e\cdot s)}}$, which consists only of 0 and 1. The mask rate denotes the proportion of 0 in Mask Matrix.
$\widetilde{\mathbf{M^{\prime}}}_{\mathrm{xy}}=\begin{cases}0,&\mathrm{~{}if~{}}0\leq\mathrm{y<r\times n\times s}\\ 1,&\mathrm{~{}else}\end{cases}$ (2)
Equation 2 denotes the element values of the original Mask Matrix $\widetilde{\mathbf{M^{\prime}}}\in\mathbb{R}^{{n\times e\times s}}$. The $r$ denotes the masking rate of the corresponding expert’s input. The $x$ and $y$ denotes the location of the elements in matrix $\widetilde{\mathbf{M^{\prime}}}$, where $x\in[0,e)$ and $y\in[0,n\cdot s)$, and both of them are natural numbers. And then, the elements in $\widetilde{\mathbf{M^{\prime}}}$ are shuffled randomly, $\tilde{\mathbf{M}}=shuffle(\widetilde{\mathbf{M^{\prime}}})$. Finally, the $\tilde{\mathbf{M}}$ is reshaped to fit the dimension of Dispatch Weights $\mathbf{D}$, $\mathbf{M}=rearrange(\tilde{\mathbf{M}},e,n,s->n,(e\cdot s))$. In practice, we employ 4 experts in each block of the backbone of CLIP’s image encoder, i.e. $e=4$. Only one expert could get the complete information of all slots, while others could only get the masked slots.
@@
And then our Masked Dispatch Weights $\mathbf{D^{\prime}}$ is the Hadamard product of Dispatch Weights $\mathbf{D}$ and Mask Matrix $\mathbf{M}$. $\mathbf{D^{\prime}}_{ij}$ and $\mathbf{M}_{ij}$ denotes the element in masked dispatch weights $\mathbf{D^{\prime}}$ and Mask Matrix $\mathbf{M}$ located in the ${i}^{th}$ column and ${j}^{th}$ row. $\mathbf{X}$ is then weighted by $\mathbf{D^{\prime}}$ and we get $\tilde{\mathbf{X}}$, the input of Experts.
@@
A proper mask rate can alleviate overfitting; however, since the texture differences between live irises and those with contact lenses occupy only a small proportion of the image, the masking ratio should be carefully controlled in practice to avoid obscuring these key differences and misleading the model’s learning process. Specifically, if there is no mask rate in the model, the overfitting problem would impact the anti-spoofing result negatively; if the mask rate is too large, the performs would gain negative impact as well because of the dropout of critical difference between liveness and spoofing samples.
IV-B Cosine Distance Loss between Experts
There are two main reasons to contribute this cosine distance loss functions. One is the inherent property of MoE, and the other one is caused by the used of Mask Mechanism. The efficiency of MoE methods is mainly due to the characteristic sparseness, which means that each expert could learn different knowledge [73]. But the knowledges between each expert are not shared, which would limit the improvement of generalization brought by MoE on iris recognition models. So, if we could find a way to let experts learn the knowledge from others, it may improve the generalization of the models. Otherwise, the mask mechanism introduces gaps between the outputs of experts handling masked and unmasked features, so we employ this cosine distance loss to mitigate these gaps.
To address these problems, based on the Mask Mechanism and Soft MoE, we employ this cosine distance loss, which could calculate the distance between the output of experts, to encourage each expert to guess the unseen part and make up the masked information. Specifically, the outputs of masked-feature experts are compared with the one of unmasked-feature experts, and their complementarities are leveraged to enhance the model’s overall generalization.
$\mathbf{C}_{ij}=\frac{\exp((\mathbf{X}\boldsymbol{\Phi})_{ij})}{\sum_{{j^{\prime}=1}}^{n}\exp((\mathbf{X}\boldsymbol{\Phi})_{{ij^{\prime}}})},$ $\mathbf{Y}=\mathbf{C}\tilde{\mathbf{Y}}$ (4)
In SoftMoE, a combine-weights is defined to combine the output slots of experts, and its definition is shown in Equation 4. The combine-weights $\mathbf{C}\in\mathbb{R}^{{d\times(e\cdot s)}}$ is with the same dimention as $\mathbf{D}$, shown in Equation 1. The $\tilde{\mathbf{Y}}\in\mathbb{R}^{{(e\cdot s)}\times d}$ is the output of experts, and its computing process is shown in Equation 5.
$\tilde{\mathbf{Y}}=Experts(\tilde{\mathbf{X}})$ (5)
Compared with the original SoftMoE, the cosine distance loss is employed before the process of Equation 5. We select the concept of cosine distance as the backbone of the loss function instead of Mean Squared Error (MSE). That is due to the aim of MoE is to measure the absolute error between two tensors, which is not suitable for evaluating the similarity between the outputs of experts. Thus, we choose the cosine distance to calculate the similarity between the output of each expert and not request the same. In detail, we masked three experts; the remaining one could see all of the information and request three masked experts’ output similar to the remaining one. As shown in Algorithm 1, all the outputs of the masked experts request to calculate the cosine distance with the unmasked one.
Equation 6 denotes the calculation process.
@@ IV-B Cosine Distance Loss between Experts
$\tilde{Y^{i}_{j}}$ denotes the $j^{th}$ slot of the $i^{th}$ Expert’s output. The Cosine Distance is the mean of the cosine similarity between the $j^{th}$ slots of the two Experts. The $e$ is the number of Experts and the $s$ is the number of slots. The $cosine\_distance_{k}$ denotes the $k^{th}$ residual block of ViT-B/16, which is the backbone of our image encoder.
As shown in Equation 7, the $cosine\_distance$ is used to calculate our mmoe loss $L_{mmoe}$.
@@
Figure 6: The feature distribution of each method on Protocol-1. Our proposed method has the clearest border and regular shape.
@@
TABLE II: This table shows the evaluation results on Protocol-1, Protocol-2, and Protocol-3 (Average).
@@
TABLE III: This table shows the specific evaluation results of each sub-protocol in Protocol-3.
V Experiment
V-A Datasets and Experimental Setting
@@
In this section, we re-state the dataset and the application of three sub-protocols of the Cross-Testing Protocol, and introduce some experimental settings, containing the evaluation metrics, the device, and other details of experiments. The datasets we employed in the experiments are those involved in our proposed Cross-Testing Protocol. For more details of the Cross-Testing Protocol, Table I lists more information. Each of three sub-protocols of the Cross-Testing Protocol has been evaluated on, to respectively confirm three different performances of iris anti-spoofing models, including our proposed MMoE. Protocol-1 evaluates the average performance of models, and Protocol-2 and Protocol-3 evaluate the cross-racial adaption and cross-device adaption respectively. We choose the Average Classification Error Rate (ACER), the Overall Detection Accuracy (ACC), the Area Under the Curve (AUC), and the Equivalent Error Rate (EER) as performance evaluation metrics. The experimental details are as follows: The backbone of CLIP’s image encoder is ViT-B/16. The mask rate is set as ”0, 0.1, 0.1, 0.1.”, and the prompts are set as ”This is an image of <label>eye.” We employ NVIDIA A100 GPUs to train and test with the Adam optimizer and the original learning rate is $10^{-6}$.
V-B Performance Evaluation
To verify the effectiveness of our proposed MMoE, some powerful deep learning models are also selected to be tested. The ResNet50 [54], ViT-B/16 [55], CLIP [33], and FLIP [32] are chosen as representative models, then benchmarked them with ACER, ACC, AUC, and EER. And their classification borders in latent space are visualized as well.
The experimental result of the basic performances of all models are shown in Table II and Table III. Table II shows all the evaluation results including our proposed MMoE with CLIP and other models on Protocol-1, Protocol-2, and Protocol-3 (Average). Table III provides the specific results of each sub-protocol in Protocol-3. To compare the performance of Ours with other methods, Figure7 and Figure9 visualize their ACER, ACC, AUC, and EER. Specifically, the Figure7 shows their performance comparison on the 6 sub-protocols of Protocol-3, and the Figure9 shows their performance comparison on Protocol-1 and Protocol-2. From the comparison, our proposed MMoE could achieve the SOTA performance on nearly all evaluation metrics for the three protocols.
@@ V-B Performance Evaluation
The evaluation results show that our proposed MMoE method (which is shortened as Ours in the following) with CLIP gets the best performance on all three sub-protocols of the Cross-Testing Protocol, among those models. For Protocol-1, Ours achieves 1.23% ACER, which is a nearly 13% improvement compared to the vanilla CLIP with 1.41% ACER. It means that our MMoE method could effectively detect spoofing iris samples on all datasets in the Cross-Testing Protocol and achieve SOTA performance. For Protocol-2, Ours achieves 7.59% ACER, much better than other previous models. It proves our proposed method has a high generalization ability on cross-racial challenges. Verified the effectiveness of our contributions. For Protocol-3, Ours achieves 4.85% ACER, which has significant advantages over other previous models. It proves our proposed method also has a high generalization ability on cross-device challenges.
Furthermore, Figure 6 demonstrates the feature distribution of each method on Protocol-1. Our proposed method has the clearest border and regular shape in the latent space, demonstrating its good ability in this iris anti-spoofing field.
In summary, upon all the above evaluations, our proposed MMoE method could effectively detect the spoofing iris images and perform high generalization on both cross-racial and cross-device situations. It achieves the SOTA performance on all protocols of the Cross-Testing Protocol.
Figure 7: This figure shows the ACER, ACC, AUC, and EER performed by ResNet50, ViT-B/16, CLIP, FLIP, and Ours conducted on the 6 sub-protocols of the Protocol-3.
@@
TABLE IV: The evaluation results of our MMoE with CLIP under different sets of mask-rates. The Rate_1 has the best performance on both Protocol-1 and Protocol-2.
Figure 8: This figure demonstrates the trends of ACERs, ACCs, AUCs, and EERs in ablation experiments, where the above one shows the results of on Protocol-1, and the below one shows the results on Protocol-2.
V-C Mask Rate Analysis
To discuss the impact of different rates of the mask, we set up 4 sets of mask rates and tested them on both Protocol-1 and Protocol-2. Among those, the first two sets mask the same rate for both three experts, and one expert could see all tokens with complete information. The difference between these two sets is that the mask rate of the first set is less than the one of the second set. The second set is to confirm that the mask rate ought to be in a reasonable range, not too large. The third one masks different rates for three experts, while one expert could also see all information as similar to the first two sets. The third set is to evaluate how the rates of different experts impact the performance of the whole model. The last one is quite different, in that all experts are masked parts of tokens, which means the MMoE model could not gain the complete input. This set demonstrates the significance of the unmasked feature. The exact numbers of the mask rates in each set are shown as follows:
Rate_1: 0, 0.1, 0.1, 0.1. Rate_2: 0, 0.5, 0.5, 0.5.
Rate_3: 0, 0.15, 0.3, 0.45. Rate_4: 0.1, 0.1, 0.1, 0.1.
Figure 9: This figure shows the ACER, ACC, AUC, and EER performed by ResNet50, ViT-B/16, CLIP, FLIP, and Ours conducted on Protocol-1 and Protocol-2, where the above one shows the results on Protocol-2 and the below one shows those on Protocol-2.
According to the evaluation results shown in Table IV, the result of the first set achieves the SOTA performance. And the results of other sets are weak in different degrees. The Rate_1 has the best performance on both Protocol-1 and Protocol-2, in which the ACERs are 1.23% and 7.59% respectively. The Rate_2 is a bit worse than the Rate_1 may caused by the high mask rate would make experts lose too much information and then influence the learning space of the MMoE model. The Rate_3 with different mask rates set has the worst performance on Protocol-1. This difference may increase the gap between the learning ability of experts and then cause a negative impact on the training process. The Rate_4 has a middle performance among these mask rate sets on Protocol-1, a bit worse than the Rate_1. And is the worst one on Protocol-2. This set makes the MMoE model could not gain the complete input information, which would passively influence the learning space.
Consequently, these experimental results further confirms that the mask rate should be in a reasonable range and a complementary feature without masking is necessary.
V-D Ablation Study